OpenStudy (anonymous):

What affect does "a" have on the parent function y equals 1 divided by x

OpenStudy (anonymous):

a=???

OpenStudy (anonymous):

idek D:

OpenStudy (anonymous):

help plz!

ganeshie8 (ganeshie8):

is there "a" in the parent function \(y = \frac{1}{x}\) ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

hmm i see only "x"

OpenStudy (anonymous):

well ya but i think the a is involved like theoretically

ganeshie8 (ganeshie8):

may i ask what is "a" ?

OpenStudy (anonymous):

idk? O.o

ganeshie8 (ganeshie8):

ok. i think it is another another variable. any variable can affect a function, only if it is part of it. so i think "a" will not affect the function "y = 1/x". only "x" can affect "y". nobody else !!

OpenStudy (anonymous):

ya i think your right

ganeshie8 (ganeshie8):

if the function was like "y=1/a" then a affects the function